OpenStudy (anonymous):

Maria has saved $600 to attend computer camp. She still needs $1,200 more to attend the computer camp. What percent of the amount required for computer camp does she have? Please show all work.

OpenStudy (amistre64):

1/3

OpenStudy (amistre64):

33.33%

OpenStudy (anonymous):

What is the formula you used?

OpenStudy (amistre64):

600+1200 = 1800 600/1800 = 6/18 = 1/3 = .3333333333....

OpenStudy (amistre64):

no formula, just common sense

OpenStudy (amistre64):

600 + 600 + 600 = how much is needed; well she has 1 out of the 3 so 33.33%

OpenStudy (anonymous):

thanks for the info. much appreciated.

OpenStudy (amistre64):

yep :)